1/1

N. N. 82,

Venu paco

translated by Amri Wandel

Bonvenu paco sur vin Bonvenu paco sur vin Bonvenu paco sur vin Bonvenu paco sur vin Paco sur vin

Translation of the hebrea poem “Hevenu shalom ” by N. N. 82 into Esperanto by Amri Wandel.

Arg-998-2018 (2014-04-27 13:37:20)

http://www.poezio.net/version?poem-id=998&version-id=2018